OpenStudy (anonymous):

Using the numbers 5, 8, and 24, create a problem using no more than four operations (addition, subtraction, multiplication, division, square, square root, cube, cube root) where the solution will be an irrational number. Explain why the result of your operations is an irrational number. (

OpenStudy (anonymous):

I dont understand how you do this?

OpenStudy (anonymous):

Irrational numbers have to use either the square root or cube root, because all the "easy" operations like add, subtract, multiply, and divide leave you with either whole numbers or fractions.. both are rational.

OpenStudy (anonymous):

Kind of a weird question...

OpenStudy (anonymous):

Ya i figured tht part out i tried almost everyhting i could think of with division multiplacation and stuff

OpenStudy (anonymous):

yes a very weird question indeed

OpenStudy (anonymous):

but like would this make an irrational (5- (24/8))

OpenStudy (anonymous):

maybe \[\sqrt{8+5+24}\] ? Let's check it...

OpenStudy (anonymous):

\[\sqrt{8+5+24} = \sqrt{37}\]

OpenStudy (anonymous):

the square root of 37 is like inbetween 6 and 7 so id say its irrational isnt it?

OpenStudy (anonymous):

that's irrational... yeah, you're correct about why too.

OpenStudy (anonymous):

YES! whoot whoot

OpenStudy (anonymous):

It doesn't say you have to use all the operations... just no more than 4... so I'd use that example with your reasoning. Good work :)

OpenStudy (anonymous):

(whoot)

OpenStudy (anonymous):

thx your an awesome helper. you know you should be a teacher

OpenStudy (anonymous):

You are very kind... I have heard that from someone else too... maybe I should listen :)

OpenStudy (anonymous):

haha yah youd definetly be good at it your a lifesaver i think ive said tht to you before :D

OpenStudy (anonymous):

That's ok :) Always glad to help if I can...

OpenStudy (anonymous):

yah well your good at it anyways thanx for helping ttyl